# Clarifying the effect of mean subtraction on Dynamic Mode Decomposition

### v5

@@ Appendix C On approximating the Koopman eigen-values via DMD
is an eigenvector-eigenvalue equation. The rows of $\mathbf{\Theta}_{1}$ are the left eigenvectors of $\mathbf{T}^{*}$ with the corresponding eigenvalues on the diagonal of $\mathbf{\Lambda}$. Hence, by Eq. 27, we have:
@@
Moving onto the second statement, we can show the forward implication by starting with the definition of $\boldsymbol{\mathbf{v}}_{\mu}$:
$\mathbf{T}^{*}\boldsymbol{\mathbf{v}}_{\mu}~{}=~{}\mu~{}\boldsymbol{\mathbf{v}}_{\mu}.$
If we pre-multiply both sides with $\mathbf{\Theta}_{1}$, it is possible to apply Eq. 37:
$\mathbf{\Theta}_{1}~{}\mathbf{T}^{*}\boldsymbol{\mathbf{v}}_{\mu}~{}=~{}\mu~{}\mathbf{\Theta}_{1}~{}\boldsymbol{\mathbf{v}}_{\mu}~{}\implies~{}\mathbf{\Lambda}(\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu})~{}=~{}\mu~{}(\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu}).$
Now, $\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}$. To see this, start with $\mathbf{X}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}$ and sequentially apply Eq. 28 followed by the full column rank of $\mathbf{C}$:
$\underbrace{\mathbf{X}}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}\iff\underbrace{\mathbf{C}}\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}\iff\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}.$
Consequently, $\mu$ is an eigenvalue of $\mathbf{\Lambda}$ and hence, by Eq. 27, $\mu\in\{\,\lambda_{i}\,\}_{i=1}^{r}$.
The backward implication can be established using the duality of the left and right eigenvectors of $\mathbf{T}^{*}$. Firstly, we note that Eq. 38 gives $\mu~{}\in~{}\{\,\hat{\lambda}_{i}\,\}_{i=1}^{n}$. So, we can let $\boldsymbol{\mathbf{w}}_{\mu}$ denote the [^17] We say “the” and not “a” because companion matrices can only have a single eigenvector for an eigenvalue, regardless of its multiplicity [10]. normalized left eigenvector of $\mathbf{T}^{*}$ corresponding to the eigenvalue $\mu$.
@@
Since $\mu\in\{\,\lambda_{i}\,\}_{i=1}^{r}$, by Eq. 37, there is a row of $\mathbf{\Theta}_{1}$ that is a scaled version of $\boldsymbol{\mathbf{w}}_{\mu}^{T}$. Hence, according to Eq. 39, we have:
@@
Pre-multiplying both sides with $\mathbf{C}$ shows that the DMD mode corresponding to $\mu$, $\mathbf{X}\boldsymbol{\mathbf{v}}_{\mu}$, has non-zero norm.
Appendix D On DMD-DFT equivalence
Lemma D.1 (A vector equality that encodes $\boldsymbol{\mathbf{\mu}}\textrm{DMD}\equiv\textrm{DFT}$).
@@ Appendix D On DMD-DFT equivalence
D.1 A necessary and sufficient condition
Proof D.3 (Proof of Theorem 4.1).
The key here is to establish the identity,
(40) $\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]=-\mathcal{P}_{\mathcal{R}(\mathbf{X}_{\rm ms}^{H})}(\mathbf{1}_{n}),$
because it reduces, by Lemma D.1, the justification of Theorem 4.1 to an exercise in linear algebra. Indeed, the forthcoming analysis is underpinned by a single linear algebraic identity:
(41) $I=\mathcal{P}_{\mathcal{N}(\mathbf{X}_{\rm ms})}+\mathcal{P}_{\mathcal{R}(\mathbf{X}_{\rm ms}^{H})}.$
Consider the least-squares problem that is solved in Algorithm 2 to build the mean-subtracted DMD model $\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]$:
(42) $\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]~{}=~{}\mathbf{X}_{\rm ms}^{\dagger}~{}(\boldsymbol{\mathbf{z}}_{n+1}-\boldsymbol{\mathbf{\mu}})~{}={}$ $\arg\min_{\boldsymbol{\mathbf{c}}}~{}~{}~{}\lVert\boldsymbol{\mathbf{c}}\rVert_{2}$
@@ D.1 A necessary and sufficient condition
Here, the feasible set,
@@
is an affine subspace, which is an object that results from the set addition of a subspace and a constant vector called the offset. In representing an affine subspace, we can use any of its elements as the offset. Since Eqs. 9 and 10 taken together tell us that $-\boldsymbol{\mathbf{1}}_{n}$ is an element of Eq. 43, the latter may be rewritten as follows:
@@
We will now show that the closest point in this affine subspace to the origin is $-\mathcal{P}_{\mathcal{R}(\mathbf{X}_{\rm ms}^{H})}(\boldsymbol{\mathbf{1}}_{n})$. Consider applying Eq. 41 to $-\boldsymbol{\mathbf{1}}_{n}$:
(44) $-\boldsymbol{\mathbf{1}}_{n}~{}=~{}\mathcal{P}_{\mathcal{N}(\mathbf{X}_{\rm ms})}(-\boldsymbol{\mathbf{1}}_{n})+\mathcal{P}_{\mathcal{R}(\mathbf{X}_{\rm ms}^{H})}(-\boldsymbol{\mathbf{1}}_{n})$
If we construct
$\tilde{\boldsymbol{\mathbf{d}}}~{}:=~{}\mathcal{P}_{\mathcal{N}(\mathbf{X}_{\rm ms})}(-\boldsymbol{\mathbf{1}}_{n})+\boldsymbol{\mathbf{d}},$
then, $\boldsymbol{\mathbf{c}}$ can be re-written thus:
$\boldsymbol{\mathbf{c}}~{}=~{}\underbrace{-\mathbf{1}_{n}}\,+\,\boldsymbol{\mathbf{d}}~{}=~{}\mathcal{P}_{\mathcal{R}(\mathbf{X}_{\rm ms}^{H})}(-\boldsymbol{\mathbf{1}}_{n})+\underbrace{\mathcal{P}_{\mathcal{N}(\mathbf{X}_{\rm ms})}(-\boldsymbol{\mathbf{1}}_{n})+\boldsymbol{\mathbf{d}}}~{}=~{}\mathcal{P}_{\mathcal{R}(\mathbf{X}_{\rm ms}^{H})}(-\boldsymbol{\mathbf{1}}_{n})+\tilde{\boldsymbol{\mathbf{d}}}.$
Now, $\boldsymbol{\mathbf{d}}\in\mathcal{N}(\mathbf{X}_{\rm ms})$ ensures that $\tilde{\boldsymbol{\mathbf{d}}}$ is also in the kernel of $\mathbf{X}_{\rm ms}$. Hence, by invoking the orthogonality of $\mathcal{N}(\mathbf{X}_{\rm ms})$ and $\mathcal{R}(\mathbf{X}_{\rm ms}^{H})$, we get:
$\lVert\boldsymbol{\mathbf{c}}\rVert^{2}=\left\lVert\mathcal{P}_{\mathcal{R}(\mathbf{X}_{\rm ms}^{H})}(-\mathbf{1}_{n})\right\rVert^{2}+\left\lVert\tilde{\boldsymbol{\mathbf{d}}}\right\rVert^{2}\geq\left\lVert\mathcal{P}_{\mathcal{R}(\mathbf{X}_{\rm ms}^{H})}(-\mathbf{1}_{n})\right\rVert^{2}.$
So, every point in Eq. 43 has a norm of at-least $\left\lVert\mathcal{P}_{\mathcal{R}(\mathbf{X}_{\rm ms}^{H})}(-\mathbf{1}_{n})\right\rVert$. Using Eq. 44, we can show that $\mathcal{P}_{\mathcal{R}(\mathbf{X}_{\rm ms}^{H})}(-\mathbf{1}_{n})$ does belong to the affine subspace Eq. 43. Consequently, it is also the closest point in Eq. 43 to the origin. Thus, by Eq. 42, we find that Eq. 40 is true.
Using Eq. 41, we can now simultaneously obtain the forward and backward implications:
$\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]~{}=~{}-\mathbf{1}_{n}\implies\mathcal{P}_{\mathcal{N}(\mathbf{X}_{\rm ms})}\mathbf{1}_{n}=0\implies\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]~{}=~{}-\mathbf{1}_{n}.$
Note that this characterization of DMD-DFT equivalence does not invoke the existence of a finite dimensional Koopman invariant subspace.
D.2 Equivalence when $\mathbf{X}_{\rm ms}$ has linearly dependent columns
@@
Proof D.6 (Proof of Theorem 4.3).
We will begin by explicitly constructing the dictionary alluded to, and then show that it indeed exhibits DMD-DFT equivalence despite $\mathbf{X}_{\rm ms}$ having linearly dependent columns.
Constructing the counter-example:
@@
Here, we have used Eq. 26 to denote the sub-matrix formed by the first $n-1$ columns of $\mathbf{\Theta}$ as $\mathbf{\Theta}_{2}$. Now, by Definition 3.8, we can define the dictionary, $\boldsymbol{\mathbf{\psi}}$, by simply specifying the matrix $\tilde{\mathbf{C}}$:
(46) $\tilde{\mathbf{C}}_{0}~{}:=~{}(\boldsymbol{\mathbf{l}}^{\perp})^{H}(\mathbf{\Lambda}-\mathbf{I})^{-1},\quad\tilde{\mathbf{C}}~{}=~{}\tilde{\mathbf{C}}_{0}\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]^{-1}.$
Validation:
@@
Our objective is to show that $\mathbf{X}_{\rm ms}$ has linearly dependent columns and also satisfies $\mathcal{P}_{\mathcal{N}(\mathbf{X}_{\rm ms})}\boldsymbol{\mathbf{1}}_{n}=\boldsymbol{\mathbf{0}}$.
Rank defectiveness of $\mathbf{X}_{\rm ms}$ follows directly from its shape. The identities Eqs. 4, 10, and 12, when taken together, tell us that $\mathbf{X}_{\rm ms}$ is of size $m\times n$. Since $\tilde{\mathbf{C}}$ is a row vector, $m=1$ and, hence, $\mathbf{X}_{\rm ms}$ is also a row vector. The condition $n\geq 2$ means $\mathbf{X}_{\rm ms}$ has at least 2 columns and these must be linearly dependent as it only has a single row.
We prove DMD-DFT equivalence by showing that the data matrix $\mathbf{X}_{\rm ms}$ satisfies a sufficient condition. To this end, we construct a banded matrix $\mathbf{B}\in\mathbb{C}^{n\times(n-1)}$ as follows:
@@
If we can show that $\mathcal{R}(\mathbf{B})=\mathcal{N}(\mathbf{X}_{\rm ms})$, then DMD-DFT equivalence is also true:
@@
Before proceeding further, we perform some algebraic simplifications. The specific choice of $\tilde{\mathbf{C}}$ in Eq. 46, when applied alongside Eq. 23 to Eq. 24, gives,
@@
We will first show that $\mathcal{R}(\mathbf{B})\subseteq\mathcal{N}(\mathbf{X}_{\rm ms})$. Consider the product $\mathbf{X}_{\rm ms}\mathbf{B}$ in light of the columns of $\mathbf{B}$ being orthogonal to $\boldsymbol{\mathbf{1}}_{n}$.
@@
$\tilde{\mathbf{C}}_{0}\underbrace{\mathbf{\Theta}_{1}\mathbf{B}}~{}=~{}\underbrace{\tilde{\mathbf{C}}_{0}(\mathbf{\Lambda}-\mathbf{I})}\mathbf{\Theta}_{2}~{}=~{}\underbrace{(\boldsymbol{\mathbf{l}}^{\perp})^{H}\mathbf{\Theta}_{2}}~{}=~{}\boldsymbol{\mathbf{0}}\implies\mathcal{R}(\mathbf{B})\subseteq\mathcal{N}(\mathbf{X}_{\rm ms}).$
Now, $\mathcal{R}(\mathbf{B})$ is a $n-1$ dimensional subspace of $\mathbb{C}^{n}$. So, if $\mathcal{R}(\mathbf{B})$ is a strict subset of $\mathcal{N}(\mathbf{X}_{\rm ms})$, then, by the rank-nullity theorem, $\mathbf{X}_{\rm ms}=\mathbf{0}$. Therefore, proving that $\mathbf{X}_{\rm ms}\neq\mathbf{0}$ is enough to show that $\mathcal{R}(\mathbf{B})=\mathcal{N}(\mathbf{X}_{\rm ms})$. We examine the first column of $\mathbf{X}_{\rm ms}$:
@@
If we expand $\boldsymbol{\mathbf{\mu}}$ as in Eq. 47, it becomes amenable to apply Lemma D.4:
$-\underbrace{\boldsymbol{\mathbf{\mu}}}+\tilde{\mathbf{C}}_{0}\mathbf{1}_{r}{}$ $=~{}\tilde{\mathbf{C}}_{0}\underbrace{\left[\mathbf{1}_{r}-(\mathbf{\Theta}\mathbf{1}_{n+1})\left(\frac{1}{n+1}\right)\right]}$

### v6

@@ Appendix D On approximating the Koopman eigen-values via DMD
Such a reduction can be achieved by expanding both $\mathbf{X}$ and $\mathbf{Y}$ according to Eq. 40, and then discarding $\mathbf{C}$, owing to its full column rank:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{X}}\mathbf{T}^{*}-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{Y}}~{}=~{}\mathbf{0}\iff{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{C}}\left(\mathbf{\Theta}_{1}\mathbf{T}^{*}-\mathbf{\Lambda}\mathbf{\Theta}_{1}\right)~{}=~{}\mathbf{0}\iff\mathbf{\Theta}_{1}\mathbf{T}^{*}-\mathbf{\Lambda}\mathbf{\Theta}_{1}~{}=~{}\mathbf{0}.$
Since $\mathbf{\Lambda}$ is a diagonal matrix, the resulting identity can be viewed as an eigenvector-eigenvalue equation.
Consequently, the construction of $\mathbf{\Lambda}$ from the Koopman eigenvalues, $\{\,\lambda_{i}\,\}_{i=1}^{r}$, gives the desired conclusion. Firstly, observe that the rows of $\mathbf{\Theta}_{1}$ are the left eigenvectors of $\mathbf{T}^{*}$ with the corresponding eigenvalues on the diagonal of $\mathbf{\Lambda}$. Now, Eq. 39 tells us that the Koopman eigenvalues, $\{\,\lambda_{i}\,\}_{i=1}^{r}$, constitute the diagonal of $\mathbf{\Lambda}$. Additionally, Eq. 8 defines the DMD eigenvalues, $\{\,\hat{\lambda}_{i}\,\}_{i=1}^{n}$, to be the eigenvalues of $\mathbf{T}^{*}$. Combining these two observations, we have:
@@
A DMD eigenvalue with a non-zero DMD mode lies in the Koopman spectrum
Moving onto the second part of Theorem 3.4, we can prove the forward implication (paraphrased above) by observing that $\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}$. To see this, start with $\mathbf{X}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}$ and sequentially apply Eq. 40 followed by the full column rank of $\mathbf{C}$:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{X}}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}\iff{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{C}}\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}\iff\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}.$
Moreover, $\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu}$ solves the eigenvector-eigenvalue equation associated with $\mathbf{\Lambda}$, for an eigenvalue of $\mu$. This can be obtained by starting with the definition of $\boldsymbol{\mathbf{v}}_{\mu}$, pre-multiplying both sides with $\mathbf{\Theta}_{1}$ and, then, applying Eq. 49:
$\mathbf{T}^{*}\boldsymbol{\mathbf{v}}_{\mu}~{}=~{}\mu~{}\boldsymbol{\mathbf{v}}_{\mu}\implies{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{\Theta}_{1}~{}\mathbf{T}^{*}}\boldsymbol{\mathbf{v}}_{\mu}~{}=~{}\mu~{}\mathbf{\Theta}_{1}~{}\boldsymbol{\mathbf{v}}_{\mu}~{}\implies~{}\mathbf{\Lambda}(\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu})~{}=~{}\mu~{}(\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu}).$
Consequently, $\mu$ is an eigenvalue of $\mathbf{\Lambda}$ and hence, by Eq. 39, also lies in the Koopman spectrum i.e., $\mu\in\{\,\lambda_{i}\,\}_{i=1}^{r}$.
A DMD mode that corresponds to a Koopman eigenvalue must be non-zero
The backward implication can be established using the duality of the left and right eigenvectors of $\mathbf{T}^{*}$. Firstly, if $\mu$ is a Koopman eigenvalue, then, Eq. 50 tells us that $\mu$ is also a DMD eigenvalue i.e., $\mu~{}\in~{}\{\,\hat{\lambda}_{i}\,\}_{i=1}^{n}$. So, we can let $\boldsymbol{\mathbf{w}}_{\mu}$ denote the [^14] We say “the” and not “a” because companion matrices can only have a single eigenvector for an eigenvalue, regardless of its multiplicity [11]. normalized left eigenvector of $\mathbf{T}^{*}$ corresponding to the eigenvalue $\mu$:
@@ A DMD mode that corresponds to a Koopman eigenvalue must be non-zero
Since $\mu\in\{\,\lambda_{i}\,\}_{i=1}^{r}$, by Eq. 49, there is a row of $\mathbf{\Theta}_{1}$ that is a scaled version of $\boldsymbol{\mathbf{w}}_{\mu}^{T}$. Hence, by Eq. 51, we have,
@@
In light of the full column rank of $\mathbf{C}$ and Eq. 40, we see that $\mathbf{X}\boldsymbol{\mathbf{v}}_{\mu}$, the DMD mode corresponding to $\mu$, has non-zero norm.
$\mathbf{\Theta}_{1}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}\implies{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{C}\mathbf{\Theta}_{1}}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}\implies\mathbf{X}\boldsymbol{\mathbf{v}}_{\mu}~{}\neq~{}\boldsymbol{\mathbf{0}}.$
[…]
Appendix E On DMD-DFT equivalence
Lemma E.1 (A vector equality that encodes $\boldsymbol{\mathbf{\mu}}\textrm{DMD}\equiv\textrm{DFT}$).
@@ Appendix E On DMD-DFT equivalence
E.1 Proof of Theorem 4.1
According to Lemma E.1, it is sufficient to establish the following identity:
(52) $\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]~{}=~{}-\mathbf{1}_{n}~{}~{}+~{}~{}\mathcal{P}_{\mathcal{N}(\mathbf{X}_{\rm ms})}\boldsymbol{\mathbf{1}}_{n}.$
This is simply because once Eq. 52 has been proven, we only need to leverage Lemma E.1 to reach the desired conclusion:
$\mathcal{P}_{\mathcal{N}(\mathbf{X}_{\rm ms})}\boldsymbol{\mathbf{1}}_{n}=\boldsymbol{\mathbf{0}}~{}\iff~{}\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]~{}=~{}-\boldsymbol{\mathbf{1}}_{n}~{}\iff~{}\boldsymbol{\mathbf{\mu}}\textrm{DMD}\equiv\textrm{DFT}.$
So, we begin by observing that $\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]$ is the closest point to the origin from an affine subspace described by the mean-subtracted data:
(53) $\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]~{}={}$ $\arg\min_{\boldsymbol{\mathbf{c}}}~{}~{}~{}\left\lVert\boldsymbol{\mathbf{c}}\right\rVert_{2}^{2}$
@@
To see this, recall that for any matrix $\mathbf{A}$ and vector $\boldsymbol{\mathbf{b}}$ of appropriate dimensions, the expression $\mathbf{A}^{\dagger}\boldsymbol{\mathbf{b}}$ can be interpreted as the minimum norm least squares solution, when $\boldsymbol{\mathbf{b}}\in\mathcal{R}(\mathbf{A})$. If we are to apply this in the context of $\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]$, which is constructed according to Eqs. 19, 5, and 6 as follows,
(54) $\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]~{}=~{}\mathbf{X}_{\rm ms}^{\dagger}(\boldsymbol{\mathbf{z}}_{n+1}-\boldsymbol{\mathbf{\mu}}),$
then, we need only check if $\boldsymbol{\mathbf{z}}_{n+1}-\boldsymbol{\mathbf{\mu}}$ lies in the column-space of $\mathbf{X}_{\rm ms}$. This can be verified from Eqs. 10 and 11 which say that the columns of $\mathbf{Z}_{\rm ms}$ sum to zero:
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{Z}_{\rm ms}}\boldsymbol{\mathbf{1}}_{n+1}~{}=~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\sum_{j=1}^{n+1}\boldsymbol{\mathbf{z}}_{j}}-\sum_{j=1}^{n+1}\boldsymbol{\mathbf{\mu}}~{}=~{}(n+1)\boldsymbol{\mathbf{\mu}}-(n+1)\boldsymbol{\mathbf{\mu}}~{}=~{}\boldsymbol{\mathbf{0}}.$
In the context of Eq. 19, this means,
(55) $\mathbf{X}_{\rm ms}\boldsymbol{\mathbf{1}}_{n}+(\boldsymbol{\mathbf{z}}_{n+1}-\boldsymbol{\mathbf{\mu}})~{}=~{}\boldsymbol{\mathbf{0}}.$
In other words, $\boldsymbol{\mathbf{z}}_{n+1}-\boldsymbol{\mathbf{\mu}}$ lies in the range-space of $\mathbf{X}_{\rm ms}$. Hence, Eq. 54 can be re-written as the minimum norm least squares problem Eq. 53.
Re-parametrizing the affine subspace gives us the following re-phrasal of Eq. 53:
(56) $\boldsymbol{\mathbf{c}}^{*}[\mathbf{Z}_{\rm ms}]~{}=~{}-\mathbf{1}_{n}~{}~{}+~{}~{}\left(\arg\min_{\boldsymbol{\mathbf{d}}\in\mathcal{N}(\mathbf{X}_{\rm ms})}~{}~{}~{}\left\lVert\mathbf{1}_{n}-\boldsymbol{\mathbf{d}}\right\rVert_{2}^{2}\right).$
This stems from the feasible set of Eq. 53,
@@
being an affine subspace, which is an object that results from the set addition of a subspace and a constant vector called the offset. In representing an affine subspace, we can use any of its elements as the offset. Since Eq. 55 tells us that $-\boldsymbol{\mathbf{1}}_{n}$ is an element of Eq. 57, the latter may be rewritten as follows:
@@
Using this re-parametrization in conjunction with the identity,
$\left\lVert-\mathbf{1}_{n}+\boldsymbol{\mathbf{d}}\right\rVert_{2}^{2}~{}=~{}\left\lVert\mathbf{1}_{n}-\boldsymbol{\mathbf{d}}\right\rVert_{2}^{2},$
we can recast Eq. 53 as Eq. 56.
Now, the best representation of $\boldsymbol{\mathbf{1}}_{n}$ in $\mathcal{N}(\mathbf{X}_{\rm ms})$ is $\mathcal{P}_{\mathcal{N}(\mathbf{X}_{\rm ms})}\boldsymbol{\mathbf{1}}_{n}$, the orthogonal projection of $\boldsymbol{\mathbf{1}}_{n}$ onto $\mathcal{N}(\mathbf{X}_{\rm ms})$:
(58) $\mathcal{P}_{\mathcal{N}(\mathbf{X}_{\rm ms})}\boldsymbol{\mathbf{1}}_{n}~{}=~{}\arg\min_{\boldsymbol{\mathbf{d}}\in\mathcal{N}(\mathbf{X}_{\rm ms})}~{}~{}~{}\left\lVert\mathbf{1}_{n}-\boldsymbol{\mathbf{d}}\right\rVert_{2}^{2}$
Feeding Eq. 58 back to Eq. 56 gives Eq. 52.
[…]
E.2 Equivalence when $\mathbf{X}_{\rm ms}$ has linearly dependent columns
@@
E.2.1 Proof of Theorem 4.3
We will begin by explicitly constructing the dictionary alluded to and, then, show that it indeed leads to DMD-DFT equivalence despite $\mathbf{X}_{\rm ms}$ having linearly dependent columns.
Constructing the counter-example
@@
Here, we have used Eq. 38 to denote the sub-matrix formed by the first $n-1$ columns of $\mathbf{\Theta}$ as $\mathbf{\Theta}_{2}$. Now, we can define the dictionary, $\boldsymbol{\mathbf{\psi}}$, as follows:
(60) $\tilde{\mathbf{C}}_{0}~{}:=~{}(\boldsymbol{\mathbf{l}}^{\perp})^{H}(\mathbf{\Lambda}-\mathbf{I})^{-1},\quad\tilde{\mathbf{C}}~{}=~{}\tilde{\mathbf{C}}_{0}\operatorname{diag}[\boldsymbol{\mathbf{\phi}}(\boldsymbol{\mathbf{\upsilon}}_{1})]^{-1},\quad\boldsymbol{\mathbf{\psi}}~{}=~{}\tilde{\mathbf{C}}\boldsymbol{\mathbf{\phi}}.$
Validation
@@
Our objective is to show that $\mathbf{X}_{\rm ms}$ has linearly dependent columns and, simultaneously, satisfies $\mathcal{P}_{\mathcal{N}(\mathbf{X}_{\rm ms})}\boldsymbol{\mathbf{1}}_{n}=\boldsymbol{\mathbf{0}}$.
Columns of $\mathbf{X}_{\rm ms}$ are linearly dependent
Rank defectiveness of $\mathbf{X}_{\rm ms}$ follows directly from its shape. The identities Eqs. 4, 11, and 19, when taken together, tell us that $\mathbf{X}_{\rm ms}$ is of size $m\times n$. Since $\tilde{\mathbf{C}}$ is a row vector, $m=1$ and, hence, $\mathbf{X}_{\rm ms}$ is also a row vector. The condition $n\geq 2$ means $\mathbf{X}_{\rm ms}$ has at least 2 columns and these must be linearly dependent as it only has a single row.
$\boldsymbol{\mathbf{\mu}}\textrm{DMD}$ is equivalent to a DFT
We can prove DMD-DFT equivalence by showing that $\mathcal{N}(\mathbf{X}_{\rm ms})$ coincides with $\mathcal{R}(\mathbf{B})$, where $\mathbf{B}\in\mathbb{C}^{n\times(n-1)}$ is the banded matrix defined thus:
@@
If we can show that $\mathcal{R}(\mathbf{B})=\mathcal{N}(\mathbf{X}_{\rm ms})$, then, Theorem 4.1 tells us that we will have DMD-DFT equivalence:
@@
Before proceeding further, we perform some algebraic simplifications. The specific choice of $\tilde{\mathbf{C}}$ in Eq. 60, when applied alongside Eqs. 35 and 36, gives,
@@
Now, we can show that $\mathcal{R}(\mathbf{B})\subseteq\mathcal{N}(\mathbf{X}_{\rm ms})$. Consider the product $\mathbf{X}_{\rm ms}\mathbf{B}$ in light of the columns of $\mathbf{B}$ being orthogonal to $\boldsymbol{\mathbf{1}}_{n}$.
@@
$\tilde{\mathbf{C}}_{0}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{\Theta}_{1}\mathbf{B}}~{}=~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\tilde{\mathbf{C}}_{0}(\mathbf{\Lambda}-\mathbf{I})}\mathbf{\Theta}_{2}~{}=~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(\boldsymbol{\mathbf{l}}^{\perp})^{H}\mathbf{\Theta}_{2}}~{}=~{}\boldsymbol{\mathbf{0}}\implies\mathcal{R}(\mathbf{B})\subseteq\mathcal{N}(\mathbf{X}_{\rm ms}).$
Now, to show that $\mathcal{R}(\mathbf{B})=\mathcal{N}(\mathbf{X}_{\rm ms})$, we need only prove that $\mathbf{X}_{\rm ms}\neq\mathbf{0}$. To see this, observe that $\mathcal{R}(\mathbf{B})$ is a $n-1$ dimensional subspace of $\mathbb{C}^{n}$. So, if $\mathcal{R}(\mathbf{B})$ is a strict subset of $\mathcal{N}(\mathbf{X}_{\rm ms})$, then, by the rank-nullity theorem, $\mathbf{X}_{\rm ms}=\mathbf{0}$. Therefore, by contraposition, when $\mathbf{X}_{\rm ms}$ is non-zero, $\mathcal{R}(\mathbf{B})$ coincides with $\mathcal{N}(\mathbf{X}_{\rm ms})$.
Presently, the first column of $\mathbf{X}_{\rm ms}$ is non-zero. To deduce this, we can start by unpacking said column of $\mathbf{X}_{\rm ms}$:
@@
If we expand $\boldsymbol{\mathbf{\mu}}$ as in Eq. 61, it becomes amenable to apply Lemma E.4:
$-{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\boldsymbol{\mathbf{\mu}}}+\tilde{\mathbf{C}}_{0}\mathbf{1}_{r}{}$ $=~{}\tilde{\mathbf{C}}_{0}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\left[\mathbf{1}_{r}-(\mathbf{\Theta}\mathbf{1}_{n+1})\left(\frac{1}{n+1}\right)\right]}$